# Ahmet Faruk Saz,  Siheng Xiong,  and Faramarz Fekri

### v1

@@ IV Node-to-Node Semantic Communication and Deductive Inference of Hypothesis
We consider a node-to-node semantic communication setting in which a transmitter node $N_{Tx}$ communicates with a receiver node $N_{Rx}$. The receiver’s objective is to identify the most plausible hypothesis $h^{*}\in\mathcal{H}=\{h_{i}\mid i\in\mathcal{I}_{\mathcal{H}}\}$, where each hypothesis $h_{i}$ is a self-consistent FOL statement expressible as in (4) and $\mathcal{I}_{\mathcal{H}}$ is the index set over hypotheses. The true world state corresponds to an unknown constituent $C^{*}\in\mathcal{C}$, and the receiver infers $h^{*}$, the hypothesis most consistent with $C^{*}$ based on messages $\{g_{t}\}_{t=1}^{T}$ received from the transmitter over $T$ rounds.
Previously, we introduced SCLD [8], a cont-information based message selection framework. At each round $t$, the transmitter, using its local observations $\mathcal{E}_{t}$, selects the semantically most content-informative FOL message $g_{t}$ of bounded size $|g_{t}|\leq B$, aiming to reduce the receiver’s uncertainty over $\mathcal{C}$ for better deduction of $h^{*}$. The transmitter does not know anything about the hypothesis set $\mathcal{H}$. As such, in our framework, the transmitter constructs and sends the message $g_{t}$ at round $t$ as:
$g_{t}^{*}=\mathop{\mathrm{arg\,max}}_{g\in\mathcal{E}_{t}\setminus\{g_{1},\dots,g_{t-1}\}}\text{cont}\left(g,\mathcal{E}_{r}\land\bigcup_{z=1}^{t-1}g_{z}\right),\text{s.t. }|g|\leq B,$ (7)
@@
Another approach to semantic communication for hypothesis inference involves accurately conveying the underlying distribution through messages $g$. In other words, the receiver’s objective becomes replicating the sender’s intended distribution as precisely as possible. This accuracy requirement can be quantitatively defined using the Kullback–Leibler (KL) divergence between the posterior distributions at the transmitter, $p(C^{w}_{t}\mid\mathcal{E}_{t})$, and at the receiver, $p(C^{w}_{r}\mid\mathcal{E}_{r})$:
@@
The transmitter then selects $g$ to minimize this divergence.
Upon receiving $g_{t}^{*}$ at round $t$, the receiver updates its inductive logical probability distribution $\mathcal{P}_{I}$ over constituents, refining its epistemic beliefs. After $T$ rounds, the receiver selects the most probable hypothesis as:
$h^{*}=\mathop{\mathrm{arg\,max}}_{h\in\mathcal{H}}c\left(h,\bigcup_{t=1}^{T}g_{t}\right).$ (9)
To analyze our framework, we consider finite evidence $\mathcal{E}$ containing $f$ distinct attributive constituents about the SotW, and let $n_{j}$ be the number of attributive constituents corresponding to $C_{j}(x)$, i.e., the number of times $n_{j}$ that has shown up in set $\mathcal{E}$. A constituent $C^{w}$ of width $w$ is compatible with $\mathcal{E}$ if $f\leq w\leq K$, where $f$ is the width of the minimal constituent and $K=4^{|\mathbf{P}|\times|\mathbf{E}_{\mathcal{E}}|^{2}}/2$ is the total number of possible attributive constituents, where $\mathbf{E}_{\mathcal{E}}$ is the set of unique entities in $\mathcal{E}$. The posterior over $\mathcal{C}$, the set of constituents, is given by:
$\mathcal{P}_{I}=\{c(C^{w},\mathcal{E})\mid f\leq w\leq K\},$ (10)
@@
$p(C^{w}\mid\mathcal{E})=\frac{p(C^{w})p(\mathcal{E}\mid C^{w})}{\sum_{w^{\prime}}p(C^{w^{\prime}})p(\mathcal{E}\mid C^{w^{\prime}})},$ (11)
If the prior $p(C^{w})$ and likelihood $p(\mathcal{E}\mid C^{w})$ are selected as in [7], the posterior $p(C^{w}\mid\mathcal{E})$ becomes:
$p(C^{w}\mid\mathcal{E})=$ (12)
@@
After applying the approximation and focusing on the dominant terms (where many constants and exponential factors cancel out), the posterior $p(C^{w}\mid\mathcal{E})$ simplifies considerably. The key dependencies involving the counts $n_{j}$ become apparent in the simplified form:
$p(C^{w}\mid\mathcal{E})\approx$ (14)
where $\alpha\geq 1$ and $\lambda\geq 0$ are constant coefficients determining the impact of priors in singular inductive inferences and inductive generalizations [7]. Next, we formulate an optimization objective to determine an optimal method of semantic communication for hypothesis deduction.
@@ V Single Round Semantic Communication as Distribution Approximation
In formal semantics grounded in first-order logic (FOL), world representations possess a well-defined structure. This structure, as declared in (1), (2), (3), and (4), permits any FOL statement including the hypothesis to have a well-defined probability assignment. Under this framework, the semantics of a hypothesis and it’s degree of confirmation can be determined. Hence, the optimal strategy for semantic communication for hypothesis deduction can be naturally formulated as the accurate transfer of this distribution. That is, the ultimate goal is for the receiver to reconstruct the intended distribution as accurately as possible. This fidelity criterion can be formalized using the Kullback–Leibler (KL) divergence between the transmitter $p_{t}=p(C^{w}_{t}\mid\mathcal{E}_{t})$ and receiver $p_{r}=p(C^{w}_{r}\mid\mathcal{E}_{r})$ posterior distributions:
@@
where $\mathbb{H}(\cdot)$ is Shannon entropy and $\mathbb{C}(\cdot)$ is cross-entropy. Since $\mathbb{H}(p_{t})$ is constant w.r.t. the transmission, minimizing $\mathcal{L}_{\text{KL}}$ is equivalent to minimizing $\mathbb{C}(p_{t},p_{r})$, or maximizing:
@@
Here, $\{n_{j}^{\prime}\}$ are the counts of the $f$ distinct attributive constituent types, as determined by the selection of message $g$, to be transmitted to the receiver (forming $\mathcal{E}_{r}$), $B$ is the total transmission budget (e.g., total number of attributive constituents communicated), and $n_{j}$ is the available number of attributive constituent $C_{j}(x)$ at the transmitter (from $\mathcal{E}_{t}$). The transmitter distribution $p_{t}(w)=p(C^{w}_{t}\mid\mathcal{E}_{t})$ is fixed throughout communication as no new evidence is observed. The receiver distribution $p_{r}(w)=p(C^{w}_{r}\mid\mathcal{E}_{r})$ depends on the chosen $\{n_{j}^{\prime}\}$ via the approximate posterior relation derived from (14):
@@
Given the symmetry, this stationary point corresponds to the global optimum of the relaxed problem [27]. This suggests evenly distributing the transmission budget $B$ across all $f$ relevant semantic constituent types when capacities are unlimited.
@@ VII Message Selection Algorithm from Candidates
The strategies above assume the transmitter can construct the exact water-filling allocation. In practice, the transmitter might have a predefined set of $M$ candidate messages, $\mathcal{M}=\{\mathcal{A}^{(1)},\dots,\mathcal{A}^{(M)}\}$. For example, these message can be identical to the M evidences that the transmitter has observed. Each candidate message $\mathcal{A}^{(m)}=\{n_{j}^{(m)}\}_{j=1}^{f}$ represents a valid allocation of semantic constituents for a single round, satisfying $\sum_{j=1}^{f}n_{j}^{(m)}=B$ and $n_{j}^{(m)}\leq n_{j}$ (assuming capacities $n_{j}$ are known or estimated). The transmitter must select $T^{\prime}$ ($T^{\prime}\leq T$) messages from $\mathcal{M}$ for transmission over $T^{\prime}$ rounds. This selection should approximate the ideal strategies outlined in Section VI.
@@
a) Approximate Long-Term Optimization: To approximate the ideal long-term strategy, the transmitter selects a subset $\mathcal{S}\subset\mathcal{M}$ of size $|\mathcal{S}|=T^{\prime}$ such that the final cumulative allocation $\mathbf{N}^{\text{final}}=\sum_{m\in\mathcal{S}}\mathcal{A}^{(m)}$ is as close as possible to the ideal total water-filling allocation $\mathrm{WF}^{\text{final}}$ calculated with budget $T^{\prime}B$. The objective is to find the subset $\mathcal{S}$ that minimizes:
@@
b) Approximate Short-Term Optimization: To approximate the ideal per-round optimization strategy, the transmitter should select messages greedily over $T^{\prime}$ rounds. At each round $t$ ($1\leq t\leq T^{\prime}$), let the cumulative allocation sent so far be $\mathbf{N}^{(t-1)}=\sum_{\tau=1}^{t-1}\mathcal{A}^{(\sigma(\tau))}$, where $\sigma(\tau)$ is the index of the message sent in round $\tau$. The ideal target cumulative allocation after $t$ rounds, according to the ideal short-term strategy, is $\mathrm{WF}^{(t)}$, representing the water-filling allocation with total budget $tB$. The transmitter selects the next message $\mathcal{A}^{(\sigma(t))}$ from the remaining candidates in $\mathcal{M}$ (ensuring capacities are not exceeded, i.e., $N_{j}^{(t-1)}+n_{j}^{(\sigma(t))}\leq n_{j}$) that minimizes the deviation from the ideal target for the current cumulative step, e.g., minimizing $\|\mathbf{N}^{(t-1)}+\mathcal{A}^{(\sigma(t))}-\mathrm{WF}^{(t)}\|$. This involves selecting both the message and its transmission order $\sigma$. A common heuristic is to greedily choose the message at round $t$ that brings the current cumulative allocation $\mathbf{N}^{(t)}$ closest to the ideal target $\mathrm{WF}^{(t)}$.
@@ VIII Convergence Analysis
Let $\mathbf{N}^{\text{final}}=\sum_{m\in\mathcal{S}}\mathcal{A}^{(m)}$ be the final cumulative allocation achieved by selecting a subset $\mathcal{S}$ of $T^{\prime}$ messages from the candidate pool $\mathcal{M}$ using the algorithm in Section VII.b, aiming to approximate the ideal long-term strategy. Let $\mathrm{WF}^{\text{final}}$ be the ideal water-filling allocation targeting budget $T^{\prime}B$. The maximum deviation satisfies:
@@
Proof Sketch. This algorithm selects the best fixed subset $\mathcal{S}$ of size $T^{\prime}$ from the candidate pool $\mathcal{M}$ to match the final target $\mathrm{WF}^{\text{final}}$. The approximation error is primarily determined by the granularity of the candidate messages in $\mathcal{M}$ and how well any combination of $T^{\prime}$ messages can match the specific target profile. Since the optimization only considers the final state and relies on a potentially limited discrete set of candidate messages, the error bound does not necessarily decrease with $T^{\prime}$ and remains bounded by a constant factor reflecting the quality and diversity of the pool $\mathcal{M}$.
@@ IX Simulation Results and Discussion
We evaluate the proposed semantic communication strategies through simulations designed to measure both the alignment between transmitter and receiver distributions and the resulting accuracy in hypothesis deduction. All derivations and simulations assume an error-free logical channel, i.e., transmitted constituents are received without distortion. This assumption isolates the effect of semantic resource allocation from channel impairments. In practice, symbol errors and partial misdecoding are unavoidable [28], and their impact on the proposed allocation strategy remains an open question. Preliminary robustness checks under small error rates indicate graceful degradation, and a systematic analysis of noisy channels constitutes an important direction for future work.

### v2

@@ IV Node-to-Node Semantic Communication and Deductive Inference of Hypothesis
We consider a node-to-node semantic communication setting in which a transmitter node $N_{Tx}$ communicates with a receiver node $N_{Rx}$. The receiver’s objective is to identify the most plausible hypothesis $h^{*}\in\mathcal{H}=\{h_{i}\mid i\in\mathcal{I}_{\mathcal{H}}\}$, where each hypothesis $h_{i}$ is a self-consistent FOL statement expressible as in Thm.  (1) and $\mathcal{I}_{\mathcal{H}}$ is the index set over hypotheses. The true world state corresponds to an unknown constituent $C^{*}\in\mathcal{C}$, and the receiver infers $h^{*}$, the hypothesis most consistent with $C^{*}$ based on messages $\{\varphi_{t}\}_{t=1}^{T}$ received from the transmitter over $T$ rounds.
Previously, we introduced SCLD [saz2024lossy], a cont-information based message selection framework. At each round $t$, the transmitter, using its local observations $\mathcal{E}_{t}$, selects the semantically most content-informative FOL message $\varphi_{t}$ of bounded size $|\varphi_{t}|\leq B$, aiming to reduce the receiver’s uncertainty over $\mathcal{C}$ for better deduction of $h^{*}$. The transmitter does not know anything about the hypothesis set $\mathcal{H}$. As such, in our framework, the transmitter constructs and sends the message $\varphi_{t}$ at round $t$ as:
$\varphi_{t}^{*}=\mathop{\mathrm{arg\,max}}_{\varphi\in\mathcal{E}_{t}\setminus\{\varphi_{1},\dots,\varphi_{t-1}\}}\text{cont}\left(\varphi,\mathcal{E}_{r}\land\bigwedge_{z=1}^{t-1}\varphi_{z}\right),\text{s.t. }|\varphi|\leq B,$ (6)
@@
Another approach to semantic communication for hypothesis inference involves accurately conveying the underlying distribution through messages $\varphi$. In other words, the receiver’s objective becomes replicating the sender’s intended distribution as precisely as possible. This accuracy requirement can be quantitatively defined using the Kullback–Leibler (KL) divergence between the posterior distributions at the transmitter, $p(C^{w,k}_{t}\mid\mathcal{E}_{t})$, and at the receiver, $p(C^{w,k}_{r}\mid\mathcal{E}_{r})$:
@@
The transmitter then selects $\varphi$ to minimize this divergence.
Upon receiving $\varphi_{t}^{*}$ at round $t$, the receiver updates its inductive logical probability distribution $\mathcal{P}_{I}$ over constituents, refining its epistemic beliefs. After $T$ rounds, the receiver selects the most probable hypothesis as:
$h^{*}=\mathop{\mathrm{arg\,max}}_{h\in\mathcal{H}}c\left(h,\bigwedge_{t=1}^{T}\varphi_{t}\right).$ (8)
To analyze our framework, we consider finite evidence $\mathcal{E}$ containing $f$ distinct attributive constituents about the SotW, and let $n_{j}$ be the number of attributive constituents corresponding to $C_{j}(x)$, i.e., the number of times $n_{j}$ that has shown up in set $\mathcal{E}$. A constituent $C^{w,k}$ of width $w$ is compatible with $\mathcal{E}$ if $f\leq w\leq K$, where $f$ is the width of the minimal constituent and $K$ is the total number of possible attributive constituents. The posterior over $\mathcal{C}$, the set of constituents, is given by:
$\mathcal{P}_{I}=\{c(C^{w,k},\mathcal{E})\mid f\leq w\leq K\},$ (9)
@@
$p(C^{w,k}\mid\mathcal{E})=\frac{p(C^{w,k})p(\mathcal{E}\mid C^{w,k})}{\sum_{w^{\prime}}p(C^{w^{\prime},k})p(\mathcal{E}\mid C^{w^{\prime},k})},$ (10)
If the prior $p(C^{w,k})$ and likelihood $p(\mathcal{E}\mid C^{w,k})$ are selected as in [c10], the posterior $p(C^{w,k}\mid\mathcal{E})$ becomes:
$p(C^{w,k}\mid\mathcal{E})=$ (11)
$\frac{\pi(\alpha,w\lambda/K)\prod_{j=1}^{f}\pi(n_{j},\lambda/w)}{\sum_{i=0}^{K-f}\left[\binom{K-f}{i}\pi\left(\alpha,(f+i)\lambda/K\right)\prod_{j=1}^{f}\pi\left(n_{j},\lambda/(f+i)\right)\right]}$
@@
After applying the approximation and focusing on the dominant terms (where many constants and exponential factors cancel out), the posterior $p(C^{w,k}\mid\mathcal{E})$ simplifies considerably. The key dependencies involving the counts $n_{j}$ become apparent in the simplified form:
$p(C^{w,k}\mid\mathcal{E})\approx$ (13)
$\frac{1}{\sum_{i=0}^{K-f}\binom{K-f}{i}\frac{\Gamma\left(\frac{w\lambda}{K}\right)}{\Gamma\left(\frac{(f+i)\lambda}{K}\right)}\alpha^{\frac{(f+i-w)\lambda}{K}}\left(\frac{\Gamma\left(\frac{\lambda}{w}\right)}{\Gamma\left(\frac{\lambda}{f+i}\right)}\right)^{f}\prod_{j=1}^{f}n_{j}^{\left(\frac{1}{f+i}-\frac{1}{w}\right)\lambda}}$
where $\alpha\geq 1$ and $\lambda\geq 0$ are constant coefficients determining the impact of priors in singular inductive inferences and inductive generalizations [c10]. Next, we formulate an optimization objective to determine an optimal method of semantic communication for hypothesis deduction.
@@ V Single Round Semantic Communication as Distribution Approximation
In formal semantics grounded in first-order logic (FOL), world representations possess a well-defined structure. This structure, as declared in (1), (2), (3), and (4), permits any FOL statement including the hypothesis to have a well-defined probability assignment. Under this framework, the semantics of a hypothesis and it’s degree of confirmation can be determined. Hence, the optimal strategy for semantic communication for hypothesis deduction can be naturally formulated as the accurate transfer of this distribution. That is, the ultimate goal is for the receiver to reconstruct the intended distribution as accurately as possible. This fidelity criterion can be formalized using the Kullback–Leibler (KL) divergence between the transmitter $p_{t}=p(C^{w,k}_{t}\mid\mathcal{E}_{t})$ and receiver $p_{r}=p(C^{w,k}_{r}\mid\mathcal{E}_{r})$ posterior distributions:
@@
where $\mathbb{H}(\cdot)$ is Shannon entropy and $\mathbb{C}(\cdot)$ is cross-entropy. Since $\mathbb{H}(p_{t})$ is constant w.r.t. the transmission, minimizing $\mathcal{L}_{\text{KL}}$ is equivalent to minimizing $\mathbb{C}(p_{t},p_{r})$, or (after simplifying logarithm) maximizing:
@@
Here, $\{n_{j}^{\prime}\}$ are the counts of the $f$ distinct attributive constituent types, as determined by the selection of message $g$, to be transmitted to the receiver (forming $\mathcal{E}_{r}$), $B$ is the total transmission budget (e.g., total number of attributive constituents communicated), and $n_{j}$ is the available number of attributive constituent $C_{j}(x)$ at the transmitter (from $\mathcal{E}_{t}$). The transmitter distribution $p_{t}(w)=p(C^{w,k}_{t}\mid\mathcal{E}_{t})$ is fixed throughout communication as no new evidence is observed. The receiver distribution $p_{r}(w)=p(C^{w,k}_{r}\mid\mathcal{E}_{r})$ depends on the chosen $\{n_{j}^{\prime}\}$ via the approximate posterior relation derived from (13):
@@
Given the symmetry, this stationary point corresponds to the global optimum of the relaxed problem [boyd2004convex]. This suggests evenly distributing the transmission budget $B$ across all $f$ relevant semantic constituent types when capacities are unlimited.
@@ VII Message Selection Algorithm from Candidates
The strategies above assume the transmitter can construct the exact water-filling allocation. In practice, the transmitter might have a predefined set of $M$ candidate messages, $\mathbf{M}=\{\mathcal{A}^{(1)},\dots,\mathcal{A}^{(M)}\}$. For example, these message can be identical to the M evidences that the transmitter has observed. Each candidate message $\mathcal{A}^{(m)}=\{n_{j}^{(m)}\}_{j=1}^{f}$ represents a valid allocation of semantic constituents for a single round, satisfying $\sum_{j=1}^{f}n_{j}^{(m)}=B$ and $n_{j}^{(m)}\leq n_{j}$ (assuming capacities $n_{j}$ are known or estimated). The transmitter must select $T^{\prime}$ ($T^{\prime}\leq T$) messages from $\mathbf{M}$ for transmission over $T^{\prime}$ rounds. This selection should approximate the ideal strategies outlined in Section VI.
@@
a) Approximate Long-Term Optimization: To approximate the ideal long-term strategy, the transmitter selects a subset $\mathcal{S}\subset\mathbf{M}$ of size $|\mathcal{S}|=T^{\prime}$ such that the final cumulative allocation $\mathbf{N}^{\text{final}}=\sum_{m\in\mathcal{S}}\mathcal{A}^{(m)}$ is as close as possible to the ideal total water-filling allocation $\mathrm{WF}^{\text{final}}$ calculated with budget $T^{\prime}B$. The objective is to find the subset $\mathcal{S}$ that minimizes:
@@
b) Approximate Short-Term Optimization: To approximate the ideal per-round optimization strategy, the transmitter should select messages greedily over $T^{\prime}$ rounds. At each round $t$ ($1\leq t\leq T^{\prime}$), let the cumulative allocation sent so far be $\mathbf{N}^{(t-1)}=\sum_{\tau=1}^{t-1}\mathcal{A}^{(\sigma(\tau))}$, where $\sigma(\tau)$ is the index of the message sent in round $\tau$. The ideal target cumulative allocation after $t$ rounds, according to the ideal short-term strategy, is $\mathrm{WF}^{(t)}$, representing the water-filling allocation with total budget $tB$. The transmitter selects the next message $\mathcal{A}^{(\sigma(t))}$ from the remaining candidates in $\mathbf{M}$ (ensuring capacities are not exceeded, i.e., $N_{j}^{(t-1)}+n_{j}^{(\sigma(t))}\leq n_{j}$) that minimizes the deviation from the ideal target for the current cumulative step, e.g., minimizing $\|\mathbf{N}^{(t-1)}+\mathcal{A}^{(\sigma(t))}-\mathrm{WF}^{(t)}\|$. This involves selecting both the message and its transmission order $\sigma$. A common heuristic is to greedily choose the message at round $t$ that brings the current cumulative allocation $\mathbf{N}^{(t)}$ closest to the ideal target $\mathrm{WF}^{(t)}$.
@@ VIII Convergence Analysis
Let $\mathbf{N}^{\text{final}}=\sum_{m\in\mathcal{S}}\mathcal{A}^{(m)}$ be the final cumulative allocation achieved by selecting a subset $\mathcal{S}$ of $T^{\prime}$ messages from the candidate pool $\mathbf{M}$ using the algorithm in Section VII.b, aiming to approximate the ideal long-term strategy. Let $\mathrm{WF}^{\text{final}}$ be the ideal water-filling allocation targeting budget $T^{\prime}B$. The maximum deviation satisfies:
@@
Proof Sketch. This algorithm selects the best fixed subset $\mathcal{S}$ of size $T^{\prime}$ from the candidate pool $\mathbf{M}$ to match the final target $\mathrm{WF}^{\text{final}}$. The approximation error is primarily determined by the granularity of the candidate messages in $\mathbf{M}$ and how well any combination of $T^{\prime}$ messages can match the specific target profile. Since the optimization only considers the final state and relies on a potentially limited discrete set of candidate messages, the error bound does not necessarily decrease with $T^{\prime}$ and remains bounded by a constant factor reflecting the quality and diversity of the pool $\mathbf{M}$.
@@ IX Simulation Results and Discussion
We evaluate the proposed semantic communication strategies through simulations designed to measure both the alignment between transmitter and receiver distributions and the resulting accuracy in hypothesis deduction. All derivations and simulations assume an error-free logical channel, i.e., transmitted constituents are received without distortion. This assumption isolates the effect of semantic resource allocation from channel impairments. In practice, symbol errors and partial misdecoding are unavoidable [yang2024harnessing], and their impact on the proposed allocation strategy remains an open question. Preliminary robustness checks under small error rates indicate graceful degradation, and a systematic analysis of noisy channels constitutes an important direction for future work.
@@ X Conclusion
Appendix A Model Theoretic Foundations of Language $\mathcal{L}$
Consider a language $\mathcal{L}$ whose syntax is defined over vocabulary consisting of a finite set of dyadic predicate symbols $\mathbf{P}=\{P_{i}\}_{i=1}^{k}$, a countably infinite set of variables $\mathbf{X}=\{x_{i}\}_{i\in\mathbb{N}}$, a countable set of constant symbols $\mathbf{C}=\{\top,\bot,$...\}, the logical connectives $\{\neg,\land,\lor\}$, the quantifiers $\{\forall,\exists\}$, and parentheses $()$ for disambiguation. Constant symbols and variables are called terms $\mathbf{T}=\mathbf{X}\cup\mathbf{C}$. The set of formulas of $\mathcal{L}$ is defined inductively, beginning with atomic formulas and extending via logical connectives and quantifiers. In particular, if $P\in\mathbf{P}$ and $t_{1},t_{2}\in\mathbf{T}$, then $P(t_{1},t_{2})$ is an atomic formula. If $F,G$ are formulas, then $\neg F$, $F\land G$, and $F\lor G$ are also formulas. Finally, if $F$ is a formula and $x_{i}$ is a variable, then both $(\exists x_{i})F$ and $(\forall x_{i})F$ are formulas.
Syntax of the language carry no meaning independent of interpretation. Interpretation is introduced into $\mathcal{L}$ by specifying models for $\mathcal{L}$. A model for $\mathcal{L}$ is a pair $\mathcal{M}=\langle\mathbf{E},J\rangle$, where $\mathbf{E}=\{e_{i}\}_{i\in\mathbb{N}}$ is a nonempty, countably infinite set called the domain (or universe), whose elements are referred to as entities. $J$ is an interpretation function assigning to each constant symbol $c_{i}\in\mathbf{C}$ entities $J(c_{i})\subseteq\mathbf{E}$, to each binary predicate symbol $P\in\mathbf{P}$ a binary relation $J(P)\subseteq\mathbf{E}\times\mathbf{E}$.
A valuation is a function $v:\mathbf{X}\to\mathbf{E}$ assigning elements of the domain to the variables of the language. Given a model $\mathcal{M}=\langle\mathbf{E},J\rangle$ and a valuation $v$, the satisfaction relation $\mathcal{M},v\models F$ determines whether a formula $F$ holds in $\mathcal{M}$ under the variable assignment specified by $v$.
In particular, an atomic formula $P(t_{1},t_{2})$ is satisfied in $\mathcal{M}$ under a valuation $v$ if and only if $(v(t_{1}),v(t_{2}))\in J(P)$. Similarly, $(\exists x)F$ is satisfied in $\mathcal{M}$ under $v$ if there exists $e_{i}\in\mathbf{E}$ such that $\mathcal{M},v[x_{i}\mapsto e_{i}]\models F$. Rest of the relations are defined recursively, following the structure of formulas.
The model $\mathcal{M}=\langle\mathbf{E},J\rangle$, together with a valuation $v$, determines the truth values of formulas via the satisfaction (entailment) relation $\models$, defined inductively on the structure of formulas. In particular, $\mathcal{M},v\models P(t_{1},t_{2})$ iff $(v(t_{1}),v(t_{2}))\in J(P)$; $\mathcal{M},v\models\neg F$ iff $\mathcal{M},v\not\models F$; $\mathcal{M},v\models F\lor G$ iff $\mathcal{M},v\models F$ or $\mathcal{M},v\models G$; and $\mathcal{M},v\models(\exists x_{i})F$ iff there exists some $e_{i}$ such that $\mathcal{M},v[x_{i}\mapsto e_{i}]\models F$. The remaining cases are defined analogously.
If $\mathcal{M},v\models F$, we say that $F$ is true in $\mathcal{M}$ under the valuation $v$. For a fixed model $\mathcal{M}$ and formula $F$, the truth of $F$ depends only on the values assigned by $v$ to the free variables $x_{i}$ of $F$. In particular, if $F$ contains no free variables, then either $\mathcal{M},v\models F$ holds for all valuations $v$ or for none.
Accordingly, if $\mathcal{M},v\models F$ for all valuations $v$, we write $\mathcal{M}\models F$ and say that $F$ is true in $\mathcal{M}$, or equivalently, that $\mathcal{M}$ is a model of $F$. More generally, if $R$ is a set of formulas and $\mathcal{M}\models R$ for every $R_{i}\in X$, then $\mathcal{M}$ is a model of $X$, written $\mathcal{M}\models X$. Finally, if $\mathcal{M}\models F$ holds for all models $\mathcal{M}$, we write $\models F$ and call $F$ (logically) valid.
Appendix B Distributive Normal Forms
Distributive normal forms for full first-order logic generalize the complete disjunctive normal forms (CDNF), also known as canonical DNFs or sums of minterms, from propositional and monadic first-order logic. In propositional logic, a CDNF is a disjunction of conjunctions in which every variable appears exactly once in each conjunction, either positively or negated. Each conjunction corresponds to a truth assignment that makes the formula true, so two formulas are logically equivalent if and only if they have the same CDNF. For example, over propositional variables $p$ and $q$, the formula $p\lor q$ has the CDNF $(p\land q)\ \lor\ (p\land\neg q)\ \lor\ (\neg p\land q)$. Constituents play an analogous role in first-order logic: they function as complete normal forms, providing maximally informative descriptions of possible configurations.
Appendix C Probability Assignment to Closed Sentences and Open Formulas with Valuations
It is important to distinguish clearly between a model and a valuation. A model $\mathcal{M}=\langle\mathbf{E},J\rangle$ specifies the domain of entities $\mathbf{E}$ and the interpretation of predicate symbols $J$, thereby fixing the relational structure of the world. A valuation $v:\mathbf{X}\to\mathbf{E}$, on the other hand, assigns elements of the domain $\mathbf{E}$ to the variables $\mathbf{X}$ of the language and serves to evaluate formulas relative to a given model.
In probability theory, probabilities are assigned to elements of a $\sigma$-algebra of events, i.e., to subsets of an outcome space $\Omega$ that are closed under complement and countable union. In logical probability, the outcome space is typically taken to be a set of possible states of the world, represented by models $\mathcal{M}$ (or by model–valuation pairs $(\mathcal{M},v)$). Accordingly, a closed first-order sentence $\varphi$, examples of which include $P(e_{1},e_{2}),\,\exists x_{1}P(x_{1},e_{2}),\,\exists x_{1}\forall x_{2}P(x_{1},x_{2})$ for entities $e_{1},e_{2}$ and variables $x_{1},x_{2}$, has a determinate truth value in every model, in the sense that either $\mathcal{M}\models\varphi$ or $\mathcal{M}\not\models\varphi$. Thus, it naturally determines an event
$\varphi]:=\{\,\mathcal{M}\in\Omega:\mathcal{M}\models\varphi\,\}.$
A probability measure $\mu$ on an appropriate $\sigma$-algebra over $\Omega$ therefore induces a probability assignment to closed sentences via $\Pr(\varphi):=\mu([\varphi])$, with the usual identities $\Pr(\neg\varphi)=1-\Pr(\varphi)$ and $\Pr(\varphi\lor\psi)=\Pr(\varphi)+\Pr(\psi)-\Pr(\varphi\land\psi)$.
By contrast, an open formula $\varphi(x_{1},\dots,x_{n})$, examples of which include $P(v[x_{1}\mapsto e_{1}],v[x_{2}\mapsto e_{2}])$, does not possess a truth value independently of a valuation $v:\mathbf{X}\to\mathbf{E}$ and hence does not, by itself, determine a subset of $\Omega$. Once a valuation is fixed, however, the satisfaction relation $\mathcal{M},v\models\varphi$ induces the event
$\varphi,v]:=\{\,\mathcal{M}\in\Omega:\mathcal{M},v\models\varphi\,\},$
which is semantically equivalent to the event determined by the corresponding grounded sentence $\varphi(c_{1},\dots,c_{n})$ with $c_{i}=v(x_{i})$, even though the open formula and the closed sentence remain syntactically distinct. Consequently, probabilities may be coherently assigned either to closed sentences or to open formulas relative to a valuation (or parameters), but not to open formulas as bare syntactic objects. This distinction reflects why only truth-valued formulas can serve as events satisfying Kolmogorov’s axioms.
This point becomes especially important when considering constituents. Constituents—understood as complete disjunctive normal forms of a first-order language—are purely syntactic constructions. They may contain free variables, provided these variables occur uniformly across all disjuncts. The presence of free variables does not obstruct their syntactic definition or manipulation, since normal-form construction proceeds independently of interpretation. However, the existence of such syntactic constituents does not warrant the assignment of probabilities to them in the absence of a model and, when relevant, a valuation. Open formulas or constituents containing free variables do not determine sets of models and therefore fail to define measurable events; assigning probabilities to them directly would undermine the set-theoretic foundations of probability by leaving complements, unions, and intersections undefined.
In inductive-logical frameworks, prior probabilities are therefore assigned to closed constituents. These priors depend only on structural symmetries of the language and not on particular valuations of variables, in the spirit of Carnap-style inductive logic [carnap_logical_1962]. Valuations enter only at the posterior stage, through Bayesian conditioning on evidence. Posterior probabilities $p(C^{w,k}\mid\mathcal{E})$ reflect how frequently particular models (and therefore constituents they logically entail) are supported by the observed data as is the true but unknown model $\mathcal{M}$ of the world $\mathbf{E}$. Since the domain of entities is assumed to be countably infinite, we further assume that only a finite subset of valuations is instantiated in the evidence.
Once a posterior distribution over constituents has been determined, probabilities can then be assigned to closed sentences, or to open formulas relative to specified valuations, by considering the set of models in which they are satisfied. In this way, free variables may appear at the syntactic level of constituent construction, but probabilistic interpretation ultimately applies only to closed sentences or to open formulas evaluated under an explicit valuation.